Genetic Insights Co. purchases an asset for $18,933. This asset qualifies as a seven-year recovery asset under MACRS. The seven-year fixed depreciation percentages for years 1, 2, 3, 4, 5, and 6 are 14.29%, 24.49%, 17.49%, 12.49%, 8.93%, and 8.93%, respectively. Genetic Insights has a tax rate of 30%. The asset is sold at the end of six years for $4,510.

Calculate tax paid on gain on disposal.
Answer #1

Book value as on date of sale=Cost-Accumulated Depreciation

=18,933*(1-0.1429-0.2449-0.1749-0.1249-0.0893-0.0893)

=$2533.2354

Hence since sale proceeds is greater than book value as on date of sale;

Gain on disposal=4,510-2533.2354

=1976.7646

Hence tax paid=1976.7646*30%

=$593.02938